As you know we are a premium golf services venue established by our owner and professional golfer Park Se Ri. The experience of a round of golf is enhanced by the technological features such as ball finders, GPS enabled 3D course maps, putt breaks, four-player scoring, and AI caddie system. All of this for the small fee of ₩80,000 per person per round. However, we need to improve profits.

In an effort to increase our profits we are thinking about raising prices by about 10% (to ₩ 90,000) since golf is increasing in popularity among Koreans and Chinese tourists.

Our sales department has supplied us with the following estimates of average monthly (inverse) demand for customers (in thousands of Korean Won): P(Q) = 120 - 6.67Q

The operations department has estimated that the total monthly cost (also in thousands of Korean Won) to the country club is: C(Q) = 400+40Q

Please review our current pricing strategy and let us know if a price hike for rounds of golf is appropriate or if you think there is an alternative pricing strategy that would improve profits more.
Answer #1

MC = dC(Q)/dQ = 40

Profit is maximized when MR = MC.

TR = P(Q) x Q = 120Q - 6.67Q2

MR = dTR/dQ = 120 - 13.34Q

120 - 13.34Q = 40

13.34Q = 80

Q = 6

P = 120 - (6.67 x 6) = 120 - 20 = 80

So, current price charged is the profit-maximizing price. No alternate strategies are required.